OpenStudy (anonymous):

You have two forces, 111.9 N and 69.6 N. What is the magnitude of their resultant if the first acts upward and the second acts downward? can somebody explain this to me plzz

OpenStudy (s3a):

The vector sum of the vertical forces (aka y direction) is: ΣF_y = (+119.9N + -69.6N) = +50.3 N (where the Σ means "add everything up" and the positive means upward and the negative means downward (as defined by us and most people). Basically, all this is is two forces fighting against each other and one wins and by doing the vector sum, we figure out which one wins, by how much, and in which direction the resultant force will be pointed in.

OpenStudy (anonymous):

resultant force= total force acting let up force be + and down force be negative so resultant force is just the addition of the two forces. +111.9 + (-69.6) = 111.9-69.6 resultant force= +42.3N so it is 42.3N upwards.